### v1

Abstract
We propose a minimal extension of the Standard Model (SM) that addresses both the smallness of neutrino masses and the dark matter (DM) puzzle via the inverse seesaw mechanism and an axion portal fermionic DM. This model generates light neutrino masses without requiring high energy scales, enhancing its testability in future collider experiments. An axion-like particle (ALP) connects the SM and DM sectors, yielding a distinct phenomenology and a positive contribution to the muon anomalous magnetic moment ($g-2$) anomaly. Our analysis shows that the model is consistent with constraints from neutrino oscillations, DM relic density, and indirect detection, as well as providing a solution to the $g-2$ anomaly. The TeV-scale scalar sector and ALP interactions suggest potential signals in collider experiments and astrophysical observations. This work offers a unified framework to address neutrino masses and DM, with implications for particle physics and cosmology.
1 Introduction
The Standard Model (SM) has been a cornerstone in understanding elementary particles and their interactions. Despite its successes, certain phenomena remain unexplained, notably the non-zero masses of neutrinos and the existence of dark matter (DM). The SM’s failure to accommodate these observations has driven significant efforts to extend it, leading to a range of proposed mechanisms and models.
One of the most profound issues with the SM is its prediction of massless neutrinos. However, experimental evidence from Super-Kamiokande Fukuda et al. (1998), SNO Ahmad et al. (2002), KamLAND Eguchi et al. (2003), Daya Bay An et al. (2024), RENOAhn et al. (2012), CHOOZ Apollonio et al. (1999), K2K Ahn et al. (2003), T2KAbe et al. (2011), and LSND Aguilar et al. (2001) confirms that neutrinos undergo flavor oscillations, and hence are massive. Traditional seesaw mechanisms have been proposed to explain the smallness of neutrino masses by introducing heavy right-handed Majorana neutrinos, which generate light neutrino masses suppressed by the heavy Majorana scale Weinberg (1979); Mohapatra and Senjanovic (1980); Schechter and Valle (1980); King (2004); Mohapatra and Valle (1986); Dev and Pilaftsis (2012), commonly referred to as the seesaw scale. However, the traditional seesaw mechanism usually requires a seesaw scale around $M\sim 10^{12}$ GeV, which greatly reduces the possibility to probe the model experimentally. On the other hand, the inverse seesaw model King and Luhn (2013); Deppisch et al. (2015); Gavela et al. (2009); Ibarra et al. (2010); Deppisch and Valle (2005); Dev and Mohapatra (2010); Das and Okada (2013) offers a more testable neutrino mass generation mechanism. In this scenario, sterile neutrinos with small, yet technically natural, lepton number-violation (LNV) couplings are introduced. These small LNV couplings help to lower the seesaw scale down to within reach of colliders. This approach has gained traction as a realistic and testable framework, with studies exploring its implications on lepton flavor violation (LFV) and collider phenomenology Langacker (1981); Davidson et al. (2008); Altarelli and Feruglio (2010); Abada and Lucente (2014); Deppisch et al. (2015); Alekhin et al. (2016); Hisano et al. (1996); Kuno and Okada (2001); Beacham et al. (2020).
On the other hand, the existence of DM has been firmly established through astrophysical observations, such as galaxy rotation curves Navas et al. (2024); Bertone et al. (2005) and cosmic microwave background measurements Aghanim et al. (2020); Spergel et al. (2003); Smoot et al. (1992). Various extensions of the SM have been proposed to accommodate DM, see Ref. Feng (2010); Arcadi et al. (2018); Bertone and Tait (2018); Cirelli et al. (2024); Bozorgnia et al. (2024); Balazs et al. (2024) for recent reviews. One of the most studied scenario is the Higgs portal DM Patt and Wilczek (2006), thanks to its simplicity and predictive nature. In this framework, the Higgs boson serves as a mediator between the DM and the SM sector. Phenomenology of the Higgs portal DM has been studied in Ref. Djouadi et al. (2012); Bishara et al. (2016); Arcadi et al. (2020). Due to its minimal structure, the Higgs portal model cannot accommodate neutrino masses without a further extension.
An alternative and compelling approach is the axion portal model, which introduces an axion or axion-like particle (ALP) as a mediator between the DM and the SM sectors. Initially proposed to address the strong CP problem in quantum chromodynamics, axion has gained interest in DM studies due to their weak interactions and stability Dine et al. (1981); Peccei and Quinn (1977); Weinberg (1978). ALP portal models offers unique signatures in indirect detection experiment, cosmic microwave background distortions, and low energy precision measurements, which can differentiate them from Higgs portal models Nomura and Thaler (2009); Freytsis et al. (2010); Bauer et al. (2023); Zhevlakov et al. (2022). It is also interesting to note that ALP can be connected to neutrino mass generation models as well Carvajal et al. (2015).
Another intriguing anomaly is the muon anomalous magnetic moment, $g-2$, where current experimental measurements at Fermilab Abi et al. (2021); Aguillard et al. (2023) and previous measurements at Brookhaven National Laboratory Bennett et al. (2006) exceed the SM prediction Aoyama et al. (2020). This discrepancy, standing at about $5\sigma$, suggests possible contributions from physics beyond the SM.  [^1] The recent lattice simulation Borsanyi et al. (2021) and experimental results Ignatov et al. (2024a, b) have called into question the validity of the SM prediction. Various models, including those with extended Higgs sectors Chen et al. (2021); Primulando et al. (2023); Iguro et al. (2023), sterile neutrinos Pinheiro et al. (2021); Ashry et al. (2023), or new gauge bosons Gninenko and Krasnikov (2001, 2018), attempt to account for this deviation by introducing additional contributions to $g-2$. In particular, models involving dark sectors or ALPs have been investigated for their potential to explain the muon $g-2$ anomaly Buen-Abad et al. (2021); Ganguly et al. (2022); Janish and Ramani (2020); deNiverville et al. (2018).
In this work, we build upon these developments by proposing a minimal SM extension that combines the inverse seesaw mechanism for neutrino mass generation and the ALP portal DM. Our model draws from the previous study that explored Higgs portals in the DM-neutrino mass model Pongkitivanichkul et al. (2019), aiming instead to utilize the ALP portal Fitzpatrick et al. (2023); Armando et al. (2024); Ghosh et al. (2024); Gutiérrez et al. (2021); Di Meglio (2021); Gola et al. (2022); Deniverville et al. (2021); Kirpichnikov et al. (2020); deNiverville and Lee (2019); deNiverville et al. (2018); Kaneta et al. (2017), which offers new dynamics and observables. In this work, we also attempt to explain the muon $g-2$ anomaly within the ALP portal model, which place a non-trivial constraint on the model parameter space. This approach not only addresses the neutrino mass and DM issues within a unified framework but also proposes testable signatures that could guide future experimental searches.
The paper is organised as follows. The model and particle contents are discussed in detail in section 2. In section 3, we discuss the model phenomenology and relevant constraints. Numerical results are presented in section 4. Finally, we summarise our findings and express an outlook in section 5.
2 The model
In this work, we propose an extension to the SM with an electroweak (EW) singlet complex scalar ($\Phi$), two EW singlet right-handed heavy neutrinos ($N^{i}_{R}$), two EW singlet left-handed heavy neutrinos ($S^{I}_{L}$), and a pair of fermionic dark matter ($\chi_{L}$ and $\chi_{R}$).  [^2] Scenarios with one $N_{R}$ and one $S_{L}$, or one $N_{R}$ and two $S_{L}$ lead to two vanishing light neutrino masses. Scenario with two $N_{R}$ and one $S_{L}$ is compatible with neutrino oscillation data. However, it would require more free parameters than the scenario presented in this work. In addition, the $U(1)_{PQ}$ global symmetry is imposed to stabilise the DM sector, and is assumed to be broken at a scale higher than the EW symmetry breaking scale. The relevant particle contents and their charge assignment under $SU(2)_{L}\times U(1)_{Y}\times U(1)_{PQ}$ is summarised in Tab. 1.
$SU(2)_{L}$ $U(1)_{Y}$ $U(1)_{PQ}$
$L$ $\mathbf{2}$ $-1/2$ $X_{N}$
$H$ $\mathbf{2}$ $+1/2$ 00
$N_{R}$ $1$ 00 $X_{N}$
$S_{L}$ $1$ 00 $X_{N}+1$
$\chi_{R}$ $1$ 00 $X_{\chi}$
$\chi_{L}$ $1$ 00 $X_{\chi}+1$
$\Phi$ $1$ 00 $1$
Table 1: The charge assignment under the electroweak and $U(1)_{PQ}$ groups.
2.1 The scalar sector
The scalar potential consistent with $U(1)_{PQ}$ is given by
$V(H,\Phi)=-\mu_{H}^{2}H^{\dagger}H+\lambda_{H}(H^{\dagger}H)^{2}-\mu_{\Phi}^{2}\Phi^{\dagger}\Phi+\lambda_{\Phi}(\Phi^{\dagger}\Phi)^{2}+\lambda_{H\Phi}(H^{\dagger}H)(\Phi^{\dagger}\Phi).$ (1)
The field $H$ and $\Phi$ can be expanded as
$H=\frac{1}{\sqrt{2}}\begin{pmatrix}\sqrt{2}G^{+}\\ v_{h}+h^{\prime}+iG^{0}\end{pmatrix},\qquad\Phi=\frac{1}{\sqrt{2}}(v_{\phi}+\phi^{\prime})e^{ia/f_{a}},$ (2)
where $G^{+}$ and $G^{0}$ are the would-be Goldstone boson. $v_{h}$, $v_{\phi}$ and $f_{a}$ are the vacuum expectation values (vev) of $H$ and $\Phi$ and the ALP decay constant, respectively. Since $U(1)_{PQ}$ is assumed to be broken at scale higher than the EW scale, we have $v_{\phi},f_{a}>v_{h}$. Note that we employ the non-linear parametrization for $\Phi$ which makes the ALP, $a$, appears explicitly as the phase. The fields $h^{\prime}$ and $\phi^{\prime}$ mix to form the mass eigenstates
$\begin{pmatrix}h\\ \phi\end{pmatrix}=\begin{pmatrix}\cos{\theta}&\sin{\theta}\\ -\sin{\theta}&\cos{\theta}\end{pmatrix}\begin{pmatrix}h^{\prime}\\ \phi^{\prime}\end{pmatrix},$ (3)
where
$\tan(2\theta)=\frac{\lambda_{H\Phi}v_{h}v_{\phi}}{\lambda_{\Phi}v_{\phi}^{2}-\lambda_{H}v_{h}^{2}}.$ (4)
Here we identify the field $h$ with the 125 GeV Higgs boson. The masses of $h$ and $\phi$ are given by
$m_{h,\phi}^{2}=\lambda_{H}v_{h}^{2}+\lambda_{\Phi}v_{\phi}^{2}\mp\sqrt{(\lambda_{\Phi}v_{\phi}^{2}-\lambda_{H}v_{h}^{2})^{2}+\lambda_{H\Phi}^{2}v_{h}^{2}v_{\phi}^{2}}\,.$ (5)
It is more convenient to express the quartic couplings $\lambda_{H}$, $\lambda_{\Phi}$, and $\lambda_{H\Phi}$ in terms of the vevs, the masses and the mixing angle. They are given by
$\lambda_{H}$ $=\frac{m_{h}^{2}\cos^{2}{\theta}+m_{\phi}^{2}\sin^{2}{\theta}}{v_{h}^{2}},$
$\lambda_{\Phi}$ $=\frac{m_{h}^{2}\sin^{2}{\theta}+m_{\phi}^{2}\cos^{2}{\theta}}{v_{\phi}^{2}},$
$\lambda_{H\Phi}$ $=\frac{(m_{\phi}^{2}-m_{h}^{2})\sin{\theta}\cos{\theta}}{v_{h}v_{\phi}}.$ (6)
Thus, the scalar potential can be described by three free parameters: $v_{\phi}$, $m_{\phi}$ and $\sin{\theta}$. We note that $\sin\theta$ is constrained by the 125 GeV Higgs boson measurement. The latest results from the ATLAS and CMS collaborations place an upper bound on $\sin\theta\leq 0.12$ Sirunyan et al. (2021); Aad et al. (2022).
The ALP potential, on the other hand, is generated from the non-perturbative effect, which breaks $U(1)_{PQ}$ into a shift symmetry, $a\to a+f_{a}$. This shift symmetry dictates that the ALP potential takes the form of the instanton cosine potential Adams et al. (2022)
$V(a)=\widetilde{\Lambda}_{a}^{4}\Big{[}1-\cos{\left(\frac{2\pi a}{f_{a}}\right)}\Big{]},$ (7)
where $\widetilde{\Lambda}_{a}^{4}$ is the maximum value of the potential. By identifying this shift under $U(1)_{PQ}$ transformation with the period of the ALP potential, one obtains a relation
$f_{a}=v_{\phi}.$ (8)
From now on we will use $f_{a}$ in place of $v_{\phi}$. The ALP potential also generates the ALP mass
$m_{a}^{2}=\frac{\Lambda_{a}^{4}}{f_{a}},$ (9)
where $\Lambda_{a}^{2}\equiv 2\pi\widetilde{\Lambda}_{a}^{2}$.
Below the $U(1)_{PQ}$ symmetry breaking scale, $f_{a}$, the interaction of ALP with other particles is given by the effective operators
$\mathcal{L}_{\rm ALP}=$ $\frac{C_{\psi\psi}}{f_{a}}(\partial_{\mu}a)\overline{\psi}\gamma^{\mu}\gamma_{5}\psi+\frac{g_{s}^{2}}{16\pi^{2}}C_{gg}\frac{a}{f_{a}}G_{\mu\nu}^{a}\tilde{G}^{a\mu\nu}$
$+\frac{g^{2}}{16\pi^{2}}C_{WW}\frac{a}{f_{a}}W_{\mu\nu}^{i}\tilde{W}^{i\mu\nu}+\frac{g^{\prime 2}}{16\pi^{2}}C_{BB}\frac{a}{f_{a}}B_{\mu\nu}\tilde{B}^{\mu\nu},$ (10)
where $\psi$ is the SM fermion with the ALP coupling $C_{\psi\psi}$. The couplings $g_{s}$, $g$, and $g^{\prime}$ are the strong, weak, and hypercharge gauge couplings respectively, and $C_{VV}$ is the gauge-ALP coupling. The dual field strength tensor is defined as $\tilde{F}_{\mu\nu}=\frac{1}{2}\epsilon_{\mu\nu\alpha\beta}F^{\alpha\beta}$. The fields $B$ and $W_{3}$ mix to become the photon and the $Z$-boson via
$\begin{pmatrix}Z\\ A\end{pmatrix}=\begin{pmatrix}c_{W}&-s_{W}\\ s_{W}&c_{W}&\end{pmatrix}\begin{pmatrix}W^{3}\\ B\end{pmatrix},$ (11)
where $s_{W}$ and $c_{W}$ are the sine and cosine of the Weinberg angle respectively. We also define $W^{\pm}_{\mu}=(W^{1}_{\mu}\mp iW^{2}_{\mu})/\sqrt{2}$. Then, the gauge-ALP interaction can be recast into the form
$\mathcal{L}_{\rm gauge-ALP}$ $=C_{\gamma\gamma}\frac{\alpha}{4\pi}\frac{a}{f_{a}}F\tilde{F}+C_{ZZ}\frac{\alpha}{4\pi s_{W}^{2}c_{W}^{2}}\frac{a}{f_{a}}Z\tilde{Z}$
$+\Bigg{(}C_{WW}\frac{\alpha}{4\pi s_{W}^{2}}\frac{a}{f_{a}}W\tilde{W}+C_{\gamma Z}\frac{\alpha}{4\pi s_{W}c_{W}}\frac{a}{f_{a}}A\tilde{Z}+{\rm h.c.}\Bigg{)},$ (12)
where we introduce a shorthand notation $F\tilde{F}=F_{\mu\nu}\tilde{F}^{\mu\nu}$, $\alpha=e^{2}/4\pi$ is the fine structure constant, and
$C_{\gamma\gamma}=C_{WW}+C_{BB},\qquad C_{\gamma Z}=c_{W}^{2}C_{WW}-s_{W}^{2}C_{BB},\quad C_{ZZ}=c_{W}^{4}C_{WW}+s_{W}^{4}C_{BB}.$ (13)
2.2 The neutrino sector
In this section, we derive the light neutrino mass matrix in the inverse seesaw scenario. The Yukawa interactions consistent with the global $U(1)_{PQ}$ are given by
$\mathcal{L}_{Yuk}\supset-Y^{\alpha i}\overline{L}^{\alpha}\tilde{H}N_{R}^{i}-Y^{\prime Ii}\Phi\overline{S_{L}}^{I}N_{R}^{i}+\mathrm{h.c.},$ (14)
where $\alpha$, $i$, and $I$ are generation indices, and $\tilde{H}=i\sigma_{2}H^{*}$. These interactions by themselves do not violate lepton number, and hence cannot generate neutrino mass. Thus, we also introduce the LNV Majorana mass terms
$\mathcal{L}_{Maj.}=-\frac{\mu^{ij}_{R}}{2}\overline{N_{R}^{c}}^{i}N_{R}^{j}-\frac{\mu^{IJ}_{S}}{2}\overline{S_{L}}^{I}{S_{L}^{c}}^{J}+\mathrm{h.c.},$ (15)
where $\psi^{c}=C\overline{\psi}^{T}$ is the charge conjugated field. The $\mu_{R}$ and $\mu_{S}$ terms also break $U(1)_{PQ}$ softly. However, they are expected to be small thanks to ’t Hooft technical naturalness argument. After $U(1)_{PQ}$ and EW symmetry breaking, Eq. (14) and (15) lead to the neutrino mass terms
$-\mathcal{L}_{\rm mass}^{\nu}=$ $m^{\alpha i}_{D}\overline{\nu^{\alpha}}_{L}N_{R}^{i}+M^{Ij}_{N}\overline{S_{L}}^{I}N_{R}^{j}$
$+\frac{\mu_{R}^{ij}}{2}\overline{N_{R}^{c}}^{i}{N_{R}}^{j}+\frac{\mu_{S}^{IJ}}{2}\overline{S_{L}}^{I}{S_{L}^{c}}^{J}+\mathrm{h.c.},$ (16)
where we define $m_{D}^{\alpha i}=Y^{\alpha i}v_{h}/\sqrt{2}$ and $M_{N}^{Ij}=Y^{\prime Ij}f_{a}/\sqrt{2}$. The neutrino mass terms can be written in a compact form
$-\mathcal{L}_{\rm mass}^{\nu}=$ $\frac{1}{2}\overline{\psi^{c}_{R}}\mathcal{M}_{\nu}\psi_{R}+\mathrm{h.c.},$ (17)
where
$\psi_{R}=\begin{pmatrix}\nu_{L}^{c}\\ N_{R}\\ S_{L}^{c}\end{pmatrix},\qquad\mathcal{M}_{\nu}=\begin{pmatrix}\mathbf{0}&m_{D}&\mathbf{0}\\ m^{T}_{D}&\mu_{R}&M^{T}_{N}\\ \mathbf{0}&M_{N}&\mu_{S}\end{pmatrix}.$ (18)
The neutrino mass matrix $\mathcal{M}_{\nu}$ can be brought to a block diagonal form by Dev and Pilaftsis (2012)
$V^{T}\mathcal{M}_{\nu}V=\begin{pmatrix}(M_{\rm light})_{3\times 3}&\bf 0\\ \bf 0&(M_{\rm heavy})_{4\times 4}\end{pmatrix},$ (19)
where $V$ is a unitary matrix, $M_{\rm light}$ and $M_{\rm heavy}$ are the light neutrino and heavy neutrino mass matrices respectively. In the limit where $||\mu_{R}||,||\mu_{S}||\ll||m_{D}||\ll||M_{N}||$  [^3] Here we define $||M||\equiv\sqrt{{\rm Tr}(M^{\dagger}M)}$, $M_{\rm light}$ can be approximated by
$M_{\rm light}\simeq-m_{D}M_{N}^{-1}\mu_{S}(M_{N}^{T})^{-1}m_{D}^{T}.$ (20)
Note that, to lowest order, the Majonana mass term $\mu_{R}$ does not contribute to $M_{\rm light}$. Notice also that in this case, $M_{\rm light}$ receives an extra suppression of order $||\mu_{S}||/||M_{N}||$ compared to the usual seesaw scenario. This makes it possible to place a seesaw scale within reach of collider experiment.
2.3 Dark matter sector
The dynamics of the DM sector is governed by the Yukawa interaction
$\mathcal{L}_{DM}=-g_{\chi}\Phi\overline{\chi_{L}}\chi_{R}+\mathrm{h.c.}.$ (21)
The $U(1)_{PQ}$ symmetry breaking generates a DM mass, $m_{\chi}=g_{\chi}f_{a}/\sqrt{2}$. Below the $U(1)_{PQ}$ breaking scale, DM interactions are given by
$\mathcal{L}_{DM}\supset$ $-\frac{m_{\chi}}{f_{a}}\sin{\theta}\;h\overline{\chi}\chi-\frac{m_{\chi}}{f_{a}}\cos{\theta}\;\phi\overline{\chi}\chi-i\frac{m_{\chi}}{f_{a}}\;a\overline{\chi}\gamma_{5}\chi$
$-i\frac{g_{\chi}\cos{\theta}}{\sqrt{2}f_{a}}a\phi\overline{\chi}\chi-i\frac{g_{\chi}\sin{\theta}}{\sqrt{2}f_{a}}ah\overline{\chi}\chi,$ (22)
where $\theta$ is the neutral scalars mixing angle defined in Eq. (4).
From the above interaction, one see that the $h$, $\phi$ and $a$ can serve as a portal between the DM sector and the SM sector. However, the $h$ and $\phi$ portal to SM fermions and gauge bosons are suppressed by $\sin^{2}2\theta\lesssim 0.06$ due to the mixing in the neutral scalar sector. The ALP portal, on the other hand, does not suffer from this suppression. For simplicity, we will assume $\theta=0$ in this work. Thus, DM phenomenology is mediated by the ALP portal.
3 Phenomenology
In this section, we study the phenomenology of our model. We are interested in the minimal scenario, in the sense of the least number of free parameters, which could explain 1) neutrino masses and mixing, 2) dark matter relic density, and 3) the muon $g-2$ anomaly.
3.1 Neutrino masses and mixing
In our scenario $m_{D}$ is a complex $3\times 2$ matrix, $M_{N}$ is a complex $2\times 2$ matrix and $\mu_{S}$ is a symmetric complex $2\times 2$ matrix. However, not all components of these matrices are physical. One can perform unitary field redefinitions on $N_{R}^{i}$ and $S_{L}^{I}$ to put $M_{N}$ and $\mu_{S}$ in the form
[…]
$\mu_{S}$ $\sim\begin{pmatrix}1&\alpha\gamma\\ \alpha\gamma&\beta\gamma^{2}\,e^{i\phi}\end{pmatrix},$ (24)
where $\gamma$, $\alpha$, $\beta$ and $\phi$ are real and positive. Note that our parametrization of $m_{N}$ and $\mu_{S}$ makes explicit that $\gamma$ drops out from the light neutrino mass matrix. In the above equations, we have also factored out the 1-1 components from both $m_{N}$ and $\mu_{S}$. Moreover, by redefining the phases of $L^{\alpha}$ and $N^{i}$, one can rotate away 4 phases in $m_{D}$. From our analysis, we find two minimal textures of $m_{D}$, which lead to viable neutrino masses and mixing. They are
$m_{D}^{(A)}$ $\sim\begin{pmatrix}1&0\\ y_{1}&y_{2}\\ 0&y_{3}\end{pmatrix},$ (25)
$m_{D}^{(B)}$ $\sim\begin{pmatrix}1&0\\ 0&y_{1}\\ y_{2}&y_{3}\end{pmatrix},$ (26)
where $y_{i}>0$. Again, we have factored out the 1-1 component of $m_{D}$. Finally, the light neutrino mass matrix are given by
$M_{\rm light}^{(A)}$ $=\kappa\begin{pmatrix}1&&y_{1}+\alpha\,y_{2}&&\alpha\,y_{3}\\ y_{1}+\alpha\,y_{2}&&y_{1}^{2}+2\alpha\,y_{1}y_{2}+\beta e^{i\phi}y_{2}^{2}&&\alpha\,y_{1}y_{3}+\beta e^{i\phi}y_{2}y_{3}\\ \alpha\,y_{3}&&\alpha\,y_{1}y_{3}+\beta e^{i\phi}y_{2}y_{3}&&\beta e^{i\phi}y_{3}^{2}\end{pmatrix},$ (27)
$M_{\rm light}^{(B)}$ $=\kappa\begin{pmatrix}1&&\alpha\,y_{1}&&y_{2}+\alpha\,y_{3}\\ \alpha\,y_{1}&&\beta e^{i\phi}y_{1}^{2}&&\alpha\,y_{1}y_{2}+\beta e^{i\phi}y_{1}y_{3}\\ y_{2}+\alpha\,y_{3}&&\alpha\,y_{1}y_{2}+\beta e^{i\phi}y_{1}y_{3}&&y_{2}^{2}+2\alpha\,y_{2}y_{3}+\beta e^{i\phi}y_{3}^{2}\end{pmatrix},$ (28)
where $\kappa\sim v^{2}||\mu_{S}||/f_{a}^{2}$ is the overall scale of neutrino masses. Thus, in our scenario, $M_{\rm light}$ depends on 7 real parameters: $\kappa$, $y_{1,2,3}$, $a$, $b$, and $\phi$.
Parameters Normal Ordering Inverted Ordering
$\sin^{2}{\theta_{12}}$ $0.308^{+0.037}_{-0.033}$ $0.308^{+0.037}_{-0.033}$
$\sin^{2}{\theta_{13}}$ $0.02215^{+0.00173}_{-0.00185}$ $0.02231^{+0.00178}_{-0.00171}$
$\sin^{2}{\theta_{23}}$ $0.470^{+0.115}_{-0.035}$ $0.550^{+0.034}_{-0.110}$
$m^{2}_{\rm solar}/10^{-5}$ eV-2 $7.49^{+0.56}_{-0.57}$ $7.49^{+0.56}_{-0.57}$
$m_{\rm atm}^{2}/10^{-3}$ eV-2 $2.513^{+0.065}_{-0.062}$ $2.484^{+0.063}_{-0.063}$
$\delta_{\rm CP}/\degree$ $212^{+152}_{-\phantom{1}88}$ $274^{+61}_{-73}$
Table 2: The best fit values of neutrino oscillation parameters and their $3\sigma$ ranges as determined from the global neutrino data Esteban et al. (2024). Here $m^{2}_{\rm solar}=m_{2}^{2}-m_{1}^{2}$ and $m_{\rm atm}^{2}=m_{3}^{2}-m_{1}^{2}$ ($m_{2}^{2}-m_{3}^{2}$) for normal ordering (inverted ordering).
The light neutrino mass matrix can be diagonalized by a unitary matrix $U$,
$U^{T}M_{\rm light}U=\text{diag.}(m_{1},m_{2},m_{3}),$ (29)
where $m_{i}>0$ are the mass eigenvalues. The above equation defines the Pontecorvo–Maki –Nakagawa–Sakata (PMNS) matrix. In general, the PMNS matrix contains three mixing angles, one Dirac phase, and two Majorana phases.
Once the neutrino mass matrix $M_{\rm light}$ has been diagonalized, the three mixing angles can be determined by
$s_{13}^{2}=|U_{e3}|^{2},\quad s_{12}^{2}=\frac{|U_{e2}|^{2}}{1-|U_{e3}|^{2}},\quad s_{23}^{2}=\frac{|U_{\mu 3}|^{2}}{1-|U_{e3}|^{2}},$ (30)
where $s_{ij}=\sin\theta_{ij}$ is the sine of the mixing angles. The Dirac phase, $\delta_{\rm CP}$, is encoded in the Jarsklog invariant
$J=\text{Im}\left(U_{\mu 3}^{\phantom{*}}U_{e3}^{*}U_{e2}^{\phantom{*}}U_{\mu 2}^{*}\right).$ (31)
The remaining two Majorana phases, in principle, can be deduced from the invariants Im[$(U_{e2}^{\phantom{*}}U_{e1}^{*})^{2}$] and Im[$(U_{e3}^{\phantom{*}}U_{e1}^{*})^{2}$enkins and Manohar (2008). However, both Majorana phases have never been determined experimentally. The values of neutrino masses, mixing angles, and the Dirac phase, as determined from the global analysis of neutrino data, are collected in Tab. 2.
An interesting feature of our set up is that the light neutrino mass matrix is rank 2. As a result, the lightest neutrino mass is vanishing. Hence, the other two non vanishing neutrino masses depend only on $\Delta m^{2}_{\rm sol}$ and $\Delta m^{2}_{\rm atm}$. In particular, the sum of neutrino masses is given by $\sum_{i}m_{i}=\sqrt{\Delta m^{2}_{\rm sol}}+\sqrt{\Delta m^{2}_{\rm atm}}$ and $\sqrt{\Delta m^{2}_{\rm atm}-\Delta m^{2}_{\rm sol}}+\sqrt{\Delta m^{2}_{\rm atm}}$ for the NO and IO scenarios respectively. The sum of neutrino masses is constrained by cosmological measurement. Combining the Planck 2018 cosmic microwave background data and the baryonic acoustic oscillation measurement result in an upper bound $\sum_{i}m_{i}\leq 130$ meV at 95% confidence level (CL) Aghanim et al. (2020). In addition to the neutrino masses sum, there are two effective masses relevant for laboratory based neutrino experiments. First, the effective mass $m_{\nu_{e}}^{\rm eff}=\sqrt{\sum_{i}|U_{ei}|^{2}m_{i}^{2}}$ determines the endpoint of beta decay spectrum. The latest measurement published by KATRIN has put an upper bound of $m_{\nu_{e}}^{\rm eff}\leq 450$ meV at 90% CL Aker et al. (2024). Second the rate for neutrinoless double beta ($0\nu\beta\beta$) decay depends on the effective mass $m_{ee}=|\sum_{i}U_{ei}^{2}m_{i}|$. The strongest upper bound on $m_{ee}$ is provided by the KamLAND-Zen experiment with $m_{ee}\leq 36-156$ meV at 90% CL Abe et al. (2023a).
3.2 Dark matter phenomonology
In addition to generating neutrino masses, our model also provides a DM candidate which is assumed to be a fermion, $\chi$. In the absence of the neutral scalar mixing, $\theta$, DM interaction with the SM sector is mediated by the ALP, see Eq. (22). Moreover, in the spirit of minimality, the ALP is taken to couple only to muons and photons to explain the muon $g-2$ anomaly, see Sec. 3.3. As a result, assuming $\phi$ and $a$ are heavier than $\chi$, there are only two DM annihilation channels
$\sigma_{\gamma\gamma}v_{\mathrm{rel}}$ $\simeq\frac{1}{4\pi^{3}}\frac{\alpha^{2}C_{\gamma\gamma}^{2}m_{\chi}^{6}}{(m_{a}^{2}-4m_{\chi}^{2})^{2}f_{a}^{4}},$ (32)
$\sigma_{\mu\bar{\mu}}v_{\rm rel}$ $\simeq\frac{8}{\pi}\frac{C_{\mu\mu}^{2}m_{\mu}^{2}m_{\chi}^{4}}{(m_{a}^{2}-4m_{\chi}^{2})^{2}f_{a}^{4}}\sqrt{1-\frac{4m_{\mu}^{2}}{m_{\chi}^{2}}},$ (33)
where $m_{\mu}$ is the mass of the muon. Note that the $\mu\bar{\mu}$ annihilation channel is parametrically suppressed by
$\frac{\sigma_{\mu\bar{\mu}}}{\sigma_{\gamma\gamma}}\sim\frac{\pi^{2}}{\alpha^{2}}\frac{C_{\mu\mu}^{2}}{C_{\gamma\gamma}^{2}}\frac{m_{\mu}^{2}}{m_{\chi}^{2}}.$ (34)
Thus, for $m_{\chi}\gtrsim 1$ TeV, one expects the $\mu\bar{\mu}$ channel to be irrelevant unless there is a large hierarchy in $C_{\mu\mu}$ and $C_{\gamma\gamma}$.
The energy density of DM is precisely measured by the Planck satellite Aghanim et al. (2020) to be
$\Omega_{\rm DM}^{\rm Planck}h^{2}=0.12\pm 0.0012.$ (35)
Assuming DM is present in the thermal bath of the early Universe, its present energy density is related to its annihilation cross section. The value of DM energy density can be determined by solving Boltzman equation numerically. However, here we will adopt a simple freeze out approximation, which yields the thermal relic cross section $\langle\sigma v_{\rm rel}\rangle=2.6\times 10^{-26}$ cm3/s. In our numerical analysis, we take ±$\pm 0.4\times 10^{-26}$ cm3/s as the nominal error associate with the approximation. The thermal relic cross section, in turns, constrains the possible range of the ALP couplings $C_{\gamma\gamma}$ and $C_{\mu\bar{\mu}}$. In our numerical analysis, we also verify that the our results based on the freeze out approximation are consistent with micrOMEGAs Belanger et al. (2002); Alguero et al. (2024). To this end, we have implemented the model files in LanHEP Semenov (2016), and then feed them to CalCHEP Belyaev et al. (2013) and micrOMEGAs for numerical computation.
The DM annihilation cross sections in Eqs. (32) and (33) also give rise to detectable signals in gamma ray telescopes. The annihilation channel $\sigma_{\gamma\gamma}v_{\mathrm{rel}}$ leads to a line signal in gamma ray spectrum at energy equals to DM mass. Such a signal has been searched for at the FermiLAT telescope in the low DM mass range, $m_{\chi}\lesssim 1$ TeV Ajello et al. (2016). The heavier DM mass range, $m_{\chi}\leq 100$ TeV, has been searched for by H.E.S.S. Abdallah et al. (2018) and MAGIC Abe et al. (2023b).
3.3 The anomalous magnetic moment of muon
As mentioned before, this model has an ability to explain the muon $g-2$ anomaly. Currently, the combination of BNAL and FNAL run 1-3 results in the discrepancy of Aguillard (2023)
$\Delta a_{\mu}\equiv a_{\mu}^{\rm EXP}-a_{\mu}^{\rm SM}=(249\pm 48)\times 10^{-11},$ (36)
where $a_{\mu}=(g-2)/2$. The discrepancy stands at a $5\sigma$ level and invites new physics interpretation.
Figure 1: The representative diagrams for axion contribution to muon $g-2$. The left and middle panels are one-loop contribution while the right is the two-loop one.
The new positive contributions to $a_{\mu}$ can be induced by the ALP-photon and ALP-muon couplings
$\mathcal{L}\supset C_{\gamma\gamma}\frac{\alpha}{4\pi}\frac{a}{f_{a}}F_{\mu\nu}\tilde{F}^{\mu\nu}+\frac{C_{\mu\mu}}{f_{a}}(\partial_{\mu}a)\overline{\mu}\gamma^{\mu}\gamma_{5}\mu,$
through the Feynman diagrams shown in Fig. 1. The analytical result for each diagram is given by Buen-Abad et al. (2021); Ganguly et al. (2022)
[…]
$\Delta a_{\mu}^{\rm(b)}=$ $-\Bigg{(}\frac{C_{\mu\mu}}{f_{a}}\Bigg{)}\Bigg{(}\frac{C_{\gamma\gamma}}{f_{a}}\Bigg{)}\frac{\alpha}{4\pi^{3}}m_{\mu}^{2}\int_{0}^{1}dx\Bigg{[}(1-x)\Bigg{(}\log{\Big{[}\frac{4\pi f_{a}^{2}}{m_{\mu}^{2}x^{2}+m_{a}^{2}(1-x)}\Big{]}}-\frac{1}{2}\Bigg{)}$
$-\frac{3}{m_{a}^{2}/m_{\mu}^{2}}\Bigg{(}x^{2}\log{\Big{[}\frac{x^{2}+m_{a}^{2}/m_{\mu}^{2}(1-x)}{x^{2}}\Big{]}}\Bigg{)}\Bigg{]},$
[…]
It should be noted that diagrams 1 (a) and (c) generate a negative and positive contribution to $\Delta a_{\mu}$ respectively. The contribution of diagram 1 (b) could be either positive or negative depending on the relative sign of $C_{\gamma\gamma}$ and $C_{\mu\mu}$.
4 Numerical results
In this section, we perform the numerical scan on the model parameter space. In our setup, the free parameters can be divided into two groups– neutrino, and DM and muon $g-2$. We first analyze the neutrino sector which depend on 7 free parameters. The free parameters are varied within the following ranges:
$1\text{ meV }\leq\kappa\leq 1\text{ eV},\quad 10^{-3}\leq y_{i},a,b\leq 10^{3},\quad 0\leq\phi\leq 2\pi.$ (38)
We search for the region of parameter space which is consistent with the neutrino oscillation data at the $3\sigma$ level. We then deduce the corresponding values for $\sum_{i}m_{i}$, $m_{\nu_{e}}^{\rm eff}$, and $m_{ee}$.
Texture $\delta_{\rm CP}$ [] $\sum_{i}m_{i}$ [meV] $m^{\mathrm{eff}}_{\nu_{e}}$ [meV] $m_{ee}$ [meV]
A 124.16--289.95 58.09--59.21 8.64--9.00 1.16--3.04
B 124.18--289.93 58.25--59.20 8.69--9.14 1.85--4.01
Table 3: The range of $\delta_{\rm CP}$, $\sum m_{i}$, $m_{\nu_{e}}^{\mathrm{eff}}$ and $m_{ee}$ in each scenario.
We find both $m_{D}^{(A)}$ and $m_{D}^{(B)}$ can only accommodate NO neutrino masses. As a result, we have $\sum_{i}m_{i}=\sqrt{\Delta m^{2}_{\rm sol}}+\sqrt{\Delta m^{2}_{\rm atm}}$ and
$\left(m_{\nu_{e}}^{\rm eff}\right)^{2}=s_{12}^{2}c_{13}^{2}\Delta m^{2}_{\rm sol}+s_{13}^{2}\Delta m^{2}_{\rm atm}.$ (39)
Moreover, due to our parametrization of the $M_{\rm light}$, we get $m_{ee}=\kappa$. The ranges of these masses consistent with neutrino oscillation data, are shown in Tab. 3. The values of these effective masses are well below their corresponding experimental bounds.
In our framework, the parameter $\phi$ is responsible for $CP$ violation. We find a strong correlation between $\phi$ and the Dirac phase $\delta$ in both scenarios A and B, see Fig. 2 (a). We also observe an upper bound $\delta\lesssim 290^{\circ}$ in both scenarios. More interestingly, we have identified a strong correlation between $m_{ee}$ and $\delta$, see Fig. 2 (b). Their relationship can be well approximated by an empirical formula
$m_{ee}^{(\rm A)}$ $=s_{12}^{2}c_{13}^{2}\sqrt{\Delta m_{\rm sol}^{2}}+s_{13}^{2}\sqrt{\Delta m_{\rm atm}^{2}}\cos{(\delta_{\rm CP})},$ (40)
$m_{ee}^{(\rm B)}$ $=s_{12}^{2}c_{13}^{2}\sqrt{\Delta m_{\rm sol}^{2}}+s_{13}^{2}\sqrt{\Delta m_{\rm atm}^{2}}\cos{(\delta_{\rm CP}+\pi)}.$ (41)
Figure 2: The relations between the free parameter $\phi$ and $\delta$ (a), and $\delta$ and $m_{ee}$ (b) in scenarios A and B.
For the DM and muon $g-2$ sector, there are 5 free parameters: $m_{\chi}$, $m_{a}$, $f_{a}$, $C_{\gamma\gamma}$ and $C_{\mu\mu}$. Here, we identify the region of parameter space consistent with the observed DM relic abundance which can also explain the muon $g-2$ anomaly. We have also taken into account constraints from gamma ray line searches. We find that in our setup, we need a rather heavy $f_{a}$ scale, $f_{a}\gtrsim 10$ TeV. Fig. 3 shows the allowed parameter space on the $m_{a}-C_{\gamma\gamma}/f_{a}$ plane for the benchmark scenario $m_{a}=m_{\phi}=1.1m_{\chi}$ with $C_{\mu\mu}/f_{a}=-0.01$, $-0.1$, $-1$ [GeV-1] and $f_{a}=10$, 100 TeV respectively. The narrow sky blue region shows the parameter space consistent with DM relic density. The green, teal, and deep blue stripes represent the parameter space which can explain the muon $g-2$ anomaly for $C_{\mu\mu}/f_{a}=-0.01$, $-0.1$ and $-1$ [GeV-1] respectively. The dotted gold region is ruled out by the overabundance of DM. The rose, coral, and orange hatched regions are ruled out by gamma ray line searches at FermiLAT Foster et al. (2023), H.E.S.S. Abdallah et al. (2018), and Magic Abe et al. (2023b) respectively. We have also included the projected exclusion region of the upcoming Cherenkov Telescope Array (CTA) Abe et al. (2024) which is shown in the crimson hatched region.
In the case of $f_{a}=10$ TeV, we find that only $C_{\mu\mu}/f_{a}=-1$ [GeV-1] can reproduce the DM density and explain the muon $g-2$ anomaly while remaining consistent with the current gamma ray line searches. However, this scenario would be probed by the upcoming CTA. On the other hand, when $f_{a}=100$ TeV, all $C_{\mu\mu}/f_{a}=-0.01$, $-0.1$ and $-1$ can provide the solutions to $g-2$ anomaly and being consistent with DM relic density and gamma ray line constraints. Unfortunately, these scenarios lie beyond the reach of CTA.
Figure 3: The region of parameter space on the $m_{a}-C_{\gamma\gamma}/f_{a}$ plane consistent with the DM relic density (sky blue) and can explain the muon $g-2$ anomaly (green, teal and deep blue for $C_{\mu\mu}/f_{a}=-0.01$, $-0.1$ and $-1$ [GeV-1] respectively). In these plots, $m_{a}=m_{\phi}=1.1m_{\chi}$. The hatched region is ruled out by the gamma ray line search and the dotted region produces an over abundance of DM. In the top (bottom) figure, $f_{a}$ = 10 (100) TeV.
5 Conclusion and discussion
We presented a minimal extension of the SM that combines the inverse seesaw mechanism with an ALP portal DM. This model addresses both the origin of neutrino masses and the nature of DM while remaining consistent with current experimental constraints. Unlike traditional seesaw models that require high mass scales, the inverse seesaw mechanism allows for TeV-scale new physics, making it potentially accessible in future collider experiments. Additionally, the ALP serves as a mediator between the SM and DM sectors, providing a distinct phenomenology that includes a positive contribution to the muon anomalous magnetic moment ($g-2$) and compatibility with observed relic density.
From our analysis, we find two minimal scenarios for the inverse seesaw mechanism. They both lead to NO neutrinos with vanishing lightest neutrino mass. As a result, the total neutrino mass $\sum_{i}m_{i}$ and the effective masses $m_{\nu_{e}}^{eff}$ and $m_{ee}$ are directly related to $\Delta m^{2}_{\rm sol}$ and $\Delta m^{2}_{\rm atm}$. For the total neutrino mass, we find $\sum_{i}m_{i}\simeq 58$ meV. This is an order of magnitude below the current Planck constraint. However, it can be probed by the Simon Observatory whose projected sensitivity is $\sum_{i}m_{i}\leq 40$ meV at 95% CL Abitbol et al. (2019). Meanwhile, the effective mass $m_{\nu_{e}}^{eff}$ in our scenarios are around 9 meV, which is an order of magnitude below the projection $m_{\nu_{e}}^{eff}\leq 200$ meV at 90% CL of both the HOLMES and the KATRIN experiments Alpert et al. (2015); Aker et al. (2019). Moreover, we find the effective mass $m_{ee}$ is of order a few meV. This is well below the projected sensitivity of $m_{ee}\leq 13-29$ meV at 90% CL of the LEGEND experiment phase II Abgrall et al. (2017). Thus, the key signature for our neutrino sector is a signal in cosmological measurements ($\sum_{i}m_{i}$) without corresponding signals in laboratory based measurements ($m_{\nu_{e}}^{eff}$ and $m_{ee}$).
In our scenario, the DM density is determined from the annihilation $\chi\bar{\chi}\to\gamma\gamma,\,\mu\bar{\mu}$ through a s-channel ALP exchange. For sufficiently large $m_{\chi}$, the diphoton annihilation channel is expected to dominate over the dimuon channel. As a result, the parameter space is tightly constrained by the gamma ray line searches at gamma ray telescopes. In the benchmark scenario where $m_{a}=m_{\phi}=1.1m_{\chi}$, we find that DM can be a thermal relic for $m_{a}\gtrsim 80$ TeV, see Fig. 3. The ALP couplings that determine DM density also generate a positive shift in the muon $g-2$. We identify the region of parameter space in the $m_{a}-C_{\gamma\gamma}/f_{a}$ plane which could explain the muon $g-2$ anomaly for $C_{\mu\mu}/f_{a}=-0.01,-0.1,-1$ [GeV-1] respectively. Most of these parameter space regions are incompatible with DM being a thermal relic and/or the constraints from gamma ray line searches. In the case where $f_{a}=10$ TeV, only the scenario where $C_{\mu\mu}/f_{a}=-1$ [GeV-1] can satisfy both constraints for $m_{a}\sim 100$ TeV. This scenario can be probed by the upcoming CTA. In the case $f_{a}=100$ TeV, all $C_{\mu\mu}/f_{a}=-0.01,-0.1,-1$ [GeV-1] can satisfy both DM relic density and gamma ray line search constraints. Unlike the $f_{a}=10$ TeV case, these region are out of reach of the CTA.
In our setup, we need $f_{a}\gtrsim 10$ TeV and $m_{a}\gtrsim 100$ TeV to be able to explain both the DM relic density and the muon $g-2$ anomaly. This would put both scalars $a$ and $\phi$ out of reach of the LHC and possibly the planed 100 TeV collider as well. The heavy neutrinos, on the other hand, can be at the scale $f_{a}$. Thus, in principle, they could be probed at the proposed 100 TeV collider such as the Future Circular Collider Bernardi et al. (2022).
Acknowledgement
This research has received funding support from the NSRF via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation [grant number B13F670063]. CP is supported by Fundamental Fund 2567 of Khon Kaen University via the National Science Research and Innovation Fund (NSRF). The work of PU is supported in part by the Mid-Career Research Grant from the National Research Council of Thailand under contract no. N42A650378. PU also thanks the High-Energy Physics Research Unit, Chulalongkorn University for hospitality while part of this work was being completed.